OpenStudy (anonymous):

what about 4√x^12 The 4 is little!

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

\[\sqrt[4]{x^(12)} = \sqrt[4]{x^3 x^3 x^3 x^3} = x^3\]

OpenStudy (anonymous):

\[√4=x3x3x3x3−−−−−−−−−√4=x3\] good as well...